# Ashish Saxena 11    Tanvir Kaur 11    Kaushik Mondal 11

### v2

@@ 3.1 Algorithm for general graph
        $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is not $\bm{-1}$: If agent $a_{i}$ does not find $parent$ component of $a_{j}^{u}(G_{2})$ to be $-1$, then node $u$ is not the root node of the current DFS traversal of agent $a_{i}$. If $a_{i}.prt\_out$ matches with $parent$ component of $a_{j}^{u}(G_{2})$, then agent $a_{i}$ moves through $a_{i}.prt\_out$. Else if $a_{i}.prt\_out$ does not match with $parent$ component of $a_{j}^{u}(G_{2})$, then agent $a_{i}$ sets $a_{i}.state=explore$ and moves through $a_{i}.prt\_out$.
    • (B)
      If $\bm{a_{i}.success=False}:$ Two cases are possible.
      • (a)
        • (i)
        • (ii)
      • (b)
        There is no agent from $\bm{G_{1}}$ with $\bm{settled=0}$ at node $\bm{u}$: It updates $a_{i}.count_{2}=a_{i}.count_{2}+1$. If $a_{i}.count_{2}<4n^{2}$, then move through $a_{i}.prt\_out$. Else, if $a_{i}.count_{2}=4n^{2}$, then $a_{i}$ updates $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$. If $a_{i}.count_{3}<4n^{2}$, then agent $a_{i}$ sets $a_{i}.state=explore$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.skip=a_{i}.prt\_out$, $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$, and moves through $a_{i}.prt\_out$. Else, if $a_{i}.count_{3}=4n^{2}$, it divides into two groups using $Group\_divide$ procedure mentioned in the algorithm for unsettled agents of $G_{1}$.
Algorithm for unsettled agent $\bm{a_{i}}$ of at even round: At even round, each unsettled agent $a_{i}$ of $G_{1}$ or $G_{2}$ updates $a_{i}.success$ parameter. If their movement in the previous rounds is successful, agent $a_{i}$ updates their $a_{i}.success=True$. Otherwise, it updates their parameter $a_{i}.success=False$.
Algorithm for settled agents: Let $a_{i}$ be at node $u$ with $a_{i}.settled=1$. It has two parameter $a_{i}^{u}(G_{1})$ and $a_{i}^{u}(G_{2})$. At each odd round based on the communication, it computes the decision of $G_{1}$ and $G_{2}$, and it stores $G_{1}$ information in $a_{i}^{u}(G_{1})$ and $G_{2}$ information in $a_{i}^{u}(G_{2})$. In move, it does not do anything. At even round, it does not do anything.
The pseudocode of our algorithm is provided in Section 6.
3.2 Correctness and analysis of algorithm
Let $m$ be the number of edges in $G$. At the beginning of the algorithm, $n+1$ agents are co-located at a node of $G$. Therefore, all agents can learn the value of $n$. Initially, each agent $a_{i}$ is part of $G_{1}$ as $a_{i}.grp\_label=10$. Assume that the adversary does not remove an edge in the first $4m$ rounds. In this case, all agents reach the dispersion configuration because this is nothing but the dispersion algorithm using the DFS traversal of [3]. Hence, all agents reach dispersion configuration in the first $4m$ rounds.
Let in round $r<4m$, all agents in $G_{1}$ be at node $u$ and want to go through edge $e=(u,\,v)$. At the round $r$, if the adversary removes an edge $e$, then the movement of agents in $G_{1}$ becomes unsuccessful. In this case, we divide the group $G_{1}$ into two new groups $G_{1}$ and $G_{2}$. And, both the groups consider the current node $u$ as the root node for the new DFS traversal and restart the algorithm for DFS via port 0. Let $r^{\prime}\geq r$ be round when agents in $G_{1}$ and $G_{2}$ are at node $u^{\prime}$ and want to go through edge $e^{\prime}=(u^{\prime},\,v^{\prime})$ via port $p$, and their movement becomes unsuccessful due to the deletion of the edge $e^{\prime}$ by the adversary. We have the following claim.
Claim 1: If edge $e^{\prime}$ does not appear within next $16n^{2}$ odd rounds after it is deleted by the adversary in round $r^{\prime}$, then agents in $G_{2}$ get dispersed in those $16n^{2}$ rounds.
Proof
In round $r^{\prime}$, for $a_{i}\in G_{2}$, $a_{i}.state$ is either $explore$ or $backtrack$. Note that each unsettled agent in $G_{2}$ has the same state in any round. Each agent $a_{i}$ takes $4m$ odd rounds to reach node $v^{\prime}$ and tries to move through edge $(v^{\prime},\,u^{\prime})$ as justified below.
  • if $\bm{a_{i}.state=backtrack}$: If agent $a_{i}\in G_{2}$ is with $a_{i}.state=backtrack$, then it starts a new DFS considering $u^{\prime}$ as a new root node and stores $a_{i}.skip=p$. In the next $4m$ odd rounds, it runs the DFS traversal on $G-\{e^{\prime}\}$. Therefore, it reaches node $v^{\prime}$ unless all agents of $G_{2}$ get dispersed.
  • if $\bm{a_{i}.state=explore}$: In this case, it continues its current DFS traversal by setting $a_{i}.prt\_out=(p+1)$ mod $\delta_{u^{\prime}}$. There are two cases: Case (i) the node $v^{\prime}$ is already visited node in its current DFS traversal by the round $r^{\prime}$, or Case (ii) the node $v^{\prime}$ is not visited in its current DFS traversal by the round $r^{\prime}$.
    Case (i): Let $a_{i}$ has taken $t$ odd rounds to visit node $v^{\prime}$ in the current DFS before trying to traverse through edge $e^{\prime}$. After round $r^{\prime}$, agent $a_{i}$ continues the DFS traversal via $(p+1)$ mod $\delta_{u^{\prime}}$ port. It may be possible that it does not reach node $v^{\prime}$ via the remaining part of the DFS traversal because it was unable to move by port $p$. According to our algorithm, agent $a_{i}\in G_{2}$ reaches the root node (say $w$) of the current DFS within $4m-t$ odd rounds (as to complete the current agents need $4m$ odd rounds) and, if there is some unsettled agent $a_{i}\in G_{2}$ at node $w$, it starts a new DFS by increasing $a_{i}.dfs\_label$ by 1, considering $w$ as root node and takes at most $t$ rounds to reach node $v^{\prime}$. In total, agent $a_{i}\in G_{2}$ takes $4m-t+t$=$4m$ odd rounds to reach node $v^{\prime}$.
    Case (ii): Suppose, the current DFS traversal of agent $a_{i}$ has been executed for $t^{\prime}$ odd rounds. To complete the current DFS, agent $a_{i}\in G_{2}$ takes $4m-t^{\prime}$ odd rounds more. In this case, agent $a_{i}$ reaches node $v^{\prime}$ within $4m-t^{\prime}$ odd rounds.
If at least one unsettled agent in $G_{2}$ is at node $v^{\prime}$, each agent $a_{i}\in G_{2}$ tries to move edge $(v^{\prime},\,u^{\prime})$. Whenever the movement is unsuccessful each agent $a_{i}\in G_{2}$ increases $a_{i}.count_{2}$ by 1. Since $e^{\prime}$ is missing for $16n^{2}$ consecutive odd rounds from round $r^{\prime}$, therefore for each agent $a_{i}\in G_{2}$, $a_{i}.count_{2}$ becomes $4n^{2}$ as from round $r^{\prime}$ to reach node $v^{\prime}$ it takes $4m$ odd round. In total time from $r^{\prime}$ passes by is $4m+4n^{2}<16n^{2}$ odd rounds. Therefore, $a_{i}.count_{2}$ reaches $4n^{2}$. In this case, agent $a_{i}\in G_{2}$ increases $a_{i}.count_{3}$\&& $a_{i}.dfs\_label$ by 1, considers node $v^{\prime}$ the root and starts the new DFS from node $v^{\prime}$ as well as stores the outgoing port of edge $(v^{\prime},\,u^{\prime})$ as $a_{i}.skip$. In the next $4m\leq 4n^{2}$ odd rounds, each agent $a_{i}\in G_{2}$ completes the DFS traversal on $G-e^{\prime}$. It is because even if it reaches node $u^{\prime}$ and wants to go through edge $e^{\prime}$ via port $p$, it reaches as $a_{i}.state=explore$. According to our algorithm, agent $a_{i}$ changes the outgoing port by $(p+1)$ mode $\delta_{u}$ and continues its DFS traversal on $G-e^{\prime}$. The total odd rounds till dispersion of agents in $G_{2}$ is at most $12n^{2}$ odd rounds. Therefore, if edge $e^{\prime}$ does not appear in consecutive $16n^{2}$ rounds (we explain later why we are using $16n^{2}$ instead of $12n^{2}$), then the group $G_{2}$ is successfully dispersed in $G$.∎
If the edge $e^{\prime}$ reappears at some round before the wait time of the group $G_{1}$, then agents in $G_{1}$ move at least one round of its current DFS. Now we make a claim that guarantees the dispersion of agents in $G_{1}$.
Claim 2: For any unsettled agent $a_{i}\in G_{2}$, if $a_{i}.count_{3}$ matches with $4n^{2}$, then agents in $G_{1}$ are already dispersed.
Proof
Let $T$ be an odd round when edge $e^{\prime}$ reappears and agents in $G_{1}$ move at least one round on its current DFS traversal. Note that $T>r^{\prime}$. After round $T$ round onwards the following cases are possible.
  • Case (A) Agents in group $G_{1}$ and $G_{2}$ want to go through the same deleted edge: In this case, the group $G_{2}$ meets with $G_{1}$ at node $u_{1}$ and both want to go through the same edge $e_{1}=(u_{1},\,v_{1})$ in round $r_{1}$ (say corresponding port $p_{1}$ of $u_{1}$). There are two scenarios: (i) $G_{1}$ reaches node $u_{1}$ in round $r^{\prime\prime}<r_{1}$ but $G_{2}$ is not present, or (ii) $G_{2}$ reaches node $u_{1}$ in round $r^{\prime\prime\prime}<r_{1}$ but $G_{1}$ is not present in round $r^{\prime\prime\prime}$. In both cases, edge $e_{1}$ is deleted from round $r^{\prime\prime}$ or $r^{\prime\prime\prime}$ onwards (note that $r^{\prime\prime},r^{\prime\prime\prime}\geq T$). Within the next $4m$ odd rounds, the other group $G_{2}$ may come to node $u_{1}$ and move through edge $e_{1}$ (this is similar to Case (i) &\&& (ii) as discussed in Claim 1). These two situations change to the case that we have discussed in Claim 1. After round $r_{1}$, if the edge is deleted for $12n^{2}$ odd rounds consecutively, then $G_{2}$ gets dispersed. Due to this reason, we are giving waiting time $16n^{2}$ instead of $12n^{2}$ to $G_{1}$.
  • Case (B) Agents in $G_{2}$ want to go through deleted edge: In this case, agents in $G_{2}$ at node $u_{2}$ try to move via edge $e_{2}=(u_{2},\,v_{2})$ using port $p_{2}$. Agent $a_{i}\in G_{2}$ has parameter $a_{i}.count_{2}$ and it is increasing this parameter by 1 whenever its movement becomes unsuccessful. If its movement is successful, then $a_{i}.count_{2}=0$. If at some odd round, say $t^{\prime}$, $a_{i}.count_{2}$ reaches to $4n^{2}$ (it means $4n^{2}$ consecutive odd rounds edge $e^{\prime}$ is missing), then agent $a_{i}\in G_{2}$ restarts the new DFS and stores $a_{i}.skip=p_{2}$ corresponding edge $e^{\prime}$ through which it is trying to move. It is important to note that if edge $e_{2}$ is missing for consecutive $4n^{2}$ odd rounds, then two cases are possible which are as follows.
    • Case (a): While $G_{2}$ is waiting for $e_{2}$ to appear, then within $4n^{2}$ odd rounds unsettled agents in $G_{1}$ get stuck at node $u_{2}$ (or $v_{2}$). If each unsettled agent $a_{j}\in G_{1}$ reaches node $u_{2}$ and wants to go through edge $e_{2}$, then it is similar to Case (A). If each unsettled agent $a_{j}\in G_{1}$ gets stuck at node $v_{2}$, then it tries to move through edge $(v_{2},\,u_{2})$ and starts increasing $a_{j}.count_{1}$ parameter by 1 in each unsuccessful movement. When $a_{i}.count_{2}=4n^{2}$, then agent $a_{i}\in G_{2}$ restarts new DFS, stores $a_{i}.skip=p_{2}$ and increases parameter $a_{i}.count_{3}$ by one. If edge $e_{2}$ does not appear next odd $4m$ rounds (note that the parameter for agent $a_{j}\in G_{1}$ is less than $16n^{2}$), then agents in $G_{2}$ achieve the dispersion. It is because even if it reaches node $v_{2}$ and wants to go through edge $(v_{2},\,u_{2})$ via some port $p$, it reaches as $a_{i}.state=explore$. Agent $a_{i}$ changes the outgoing port by $(p+1)$ mode $\delta_{v_{2}}$ and continues its DFS traversal on $G-\{e_{2}\}$. Therefore, if edge $e_{2}$ does not appear in the next consecutive $4m$ odd rounds, then the group $G_{2}$ is successfully dispersed at the nodes of $G$. If edge $e_{2}$ reappears then agents in $G_{1}$ move at least one round of its current DFS.
It is important to note that if unsettled agent $a_{i}\in G_{2}$ increases $a_{i}.count_{3}$, then agent $a_{j}\in G_{1}$ moves at least one round of its current DFS. Therefore, whenever $a_{i}.count_{3}$ reaches $4n^{2}$, agents in $G_{2}$ can understand that agents in $G_{1}$ are dispersed as $G_{1}$ already ran its dispersion algorithm via DFS for at least $4m$ odd rounds without deviating from the path. ∎
Based on Claim 1 &\&& 2, we have the following lemma.
Lemma 2
Our algorithm correctly solves $(n+1)$-Dispersion on TVG for co-located $n+1$ agents.
Proof
Due to Claim 1 &\&& 2, we can say either $G_{1}$ or $G_{2}$ is dispersed and unsettled agents in $G_{1}\,(G_{2})$ understand whether agents in $G_{2}\,(G_{1})$ are dispersed. And, agents in $G_{1}\,($or $G_{2})$ divide into two new groups. In this way, in the end, group $G_{1}$ (and $G_{2}$) contains 1 agent. Both groups start running our algorithm, and due to Claim 1 &\&& 2, we can say that either agent in $G_{1}$ or $G_{2}$ is settled at some node. Therefore, our algorithm correctly solves $(n+1)$-Dispersion on TVG for co-located $n+1$ agents. ∎
Due to Lemma 2, we can say that one group is divided into two groups only after the other group achieves dispersion. Since we have $n+1$ agents, in this way, all agents are getting dispersed except one agent which will be part of the last remaining group.
Lemma 3
The time complexity of our algorithm is $O(n^{4}\cdot\log n)$.
Proof
In Lemma 2, we have shown that the agents achieve the dispersion. If the adversary does not remove any edge in the first $4m$ odd rounds then all agents are dispersed except one agent. In odd round $r$ (where $r<4m$), if the adversary removes the edge $e$ and the group of agents wants to go through the edge $e$ according to their DFS algorithm, then they divide into two groups. We show that in the next $32n^{4}$ odd rounds, either $G_{1}$ or $G_{2}$ gets dispersed. We have proved that whenever $a_{i}\in G_{2}$ increases $a_{i}.count_{3}$ by 1, then agents in $G_{1}$ move at least one round of its DFS or agents in $G_{1}$ get dispersed. Therefore, if agents in $G_{2}$ are not dispersed and for $a_{i}\in G_{2}$, $a_{i}.count_{3}$ reaches $4n^{2}$, then agents in $G_{2}$ can understand that agents in $G_{1}$ are dispersed. Suppose agents $G_{2}$ runs each DFS algorithm for $4m$ rounds, therefore $4(m+n^{2})\cdot 4n^{2}\leq 32n^{4}$ odd rounds are required to disperse $G_{1}$. We also have shown if $G_{1}$ waits for an edge to reappear for $16n^{2}$ rounds, then $G_{2}$ is dispersed. Therefore, after dividing into two groups, in the next $32n^{4}$ odd rounds, either $G_{1}$ or $G_{2}$ is dispersed. Therefore, either $G_{1}$ or $G_{2}$ gets dispersed in the first $32n^{4}+4m$ odd rounds. Including even rounds, we can say that after dividing into two groups, in the next $64n^{4}+8m$ rounds, either $G_{1}$ or $G_{2}$ is dispersed.
Using Lemma 2, if $G_{1}$ (or $G_{2}$) understands that $G_{2}$ (or $G_{1}$) is dispersed, then $G_{1}$ (or $G_{2}$) divides into two new groups, and both groups contain half agents of $G_{1}$ (or $G_{2}$). These two groups repeat the same procedure. The number of such group divisions is $O(\log n)$, and one of two groups is dispersed in $64n^{4}+8m$ rounds. Therefore, our algorithm takes $O(n^{4}\cdot\log n)$ rounds to solve $(n+1)$-Dispersion on TVG for co-located $n+1$ agents. ∎
Lemma 4
In our algorithm, agents require $\Theta(\log n)$ memory.
Proof
For agent $a_{i}$, $a_{i}.grp\_label$ is nothing but a binary string. Whenever we divide the group, the $a_{i}.grp\_label$ is appended by 0 or 1. The number of group divisions is $O(\log n)$. Therefore, the length of $a_{i}.grp\_label$ is $O(\log n)$, which agent $a_{i}$ can store in its $O(\log n)$ memory. According to Lemma 3, all agents achieve dispersion in $O(n^{4}\cdot\log n)$ rounds. Therefore, $a_{i}.dfs\_label$ can not be more than $O(n^{4}\cdot\log n)$ which agent $a_{i}$ can store in its $O(\log n)$ memory. Apart from these two parameters, all unsettled agents $a_{i}$ either store boolean variables or port information. Such parameters are constant as per our algorithm. To store each parameter, we need $O(\log n)$ storage. Therefore, each unsettled agent $a_{i}$ needs $O(\log n)$ memory. All settled agents maintain two parameters as per our algorithm $a_{i}^{u}(G_{1})$ and $a_{i}^{u}(G_{2})$. To store each parameter, settled agents need $O(\log n)$ memory. Therefore, each settled agent $a_{i}$ needs $O(\log n)$ memory.
Due to Lemma 1, the memory lower bound for the dispersion of $k\geq n+1$ agents is $\Omega(\log n)$, and in our algorithm agents use $O(\log n)$ memory each. Therefore, the agents in our algorithm require $\Theta(\log n)$ memory due to $k=n+1$. ∎
Now, we are ready to provide our main theorem.
Theorem 3.1
Our algorithm solves $k$-Dispersion on TVG for $k=n+1$ co-located agents in $\mathcal{G}$ in $O(n^{4}\cdot\log n)$ rounds using $\Theta(\log n)$ memory per agent.
Proof
The proof follows from Lemma 2, 3, and 4. ∎
Observation 1
All agents know the value of $n$ and $m\leq n^{2}$. If all agents want to achieve the termination, can calculate the value of $72n^{4}\cdot\lceil\log n\rceil$ and terminate after $72n^{4}\cdot\lceil\log n\rceil$ rounds.
4 Discussion
In this section, we give a discussion of how to solve $(pn+1)$-Dispersion on TVG without global communication using 0-hop visibility. We also provide an algorithm to solve $(pn)$-Dispersion on TVG with global communication and 1-hop visibility using the algorithm of [16].
4.1 $(pn+1)$-Dispersion on TVG
In this section, we demonstrate how to extend the algorithm from Section 3 to address the $(pn+1)$-Dispersion on TVGs, where $p>1$. In the dispersed configuration, each node contains $p$ agents, except one node contains $p+1$ agents. We assume that agents are aware of the parameter $n$, and $pn+1$ agents are co-located at some node $v$. In the first round, agents distribute into $n$ groups, say $g_{1}$, $g_{2}$, …, $g_{n}$. Since agents are co-located, they can compute the value of $p$ using $n$. The number of agents in each $g_{i}$, $11\leq i\leq n-$ 1, is $p$, and $g_{n}$ is $p+1$. Each agent $a_{i}$ in $g_{j}$ remembers $j$ in $a_{i}.ID_{1}$ parameter, where $1\leq j\leq n$. In the same manner, agents start executing the algorithm mentioned in Section 3. In this algorithm, whenever agents find a node without any agents, then $g_{i}$ gets settled in place of an agent. Here, the remaining unsettled $g_{i}$s divide into two groups, say $G_{1}$ and $G_{2}$, in place of dividing groups of agents into 2 groups like Section 3. In Lemma 2, we have shown one of $G_{1}$ or $G_{2}$ gets dispersed based on their parameters. When $G_{1}$ ($G_{2}$) is left with one $g_{j}$ (unsettled group), and $G_{1}$ ($G_{2}$) is sure that $G_{2}$ ($G_{1}$) has been dispersed, then they follow the following steps. W.l.o.g., let $G_{1}$ be with one $g_{i}$ at node $v$, and $G_{2}$ has been dispersed. In this case, it is important to note that 1 node is left, which has no agents. If $v$ is such a node, then dispersion has been achieved. If $v$ is not such a node, then there is some node $u$ which has no agent. In our algorithm mentioned in Section 3, there were two agents left at the end who were searching for the single empty node via different paths, one gets to that node as the adversary can not stop both at a round, and the other settled when it is sure that it the only one that is left to settle. Here, we can not use our algorithm idea because the adversary can always stop one group, i.e., the last remaining group, to find the single empty node in the graph. Below we provide the idea with correctness how the group $g_{i}$ gets settled at node $u$. There are two cases, and we discuss each case as follows.
  1. 1.
    Case $\bm{|g_{i}|=p+1}$ : The agents start running the algorithm mentioned in Section 3 in the following manner. Now $g_{i}$ executes DFS for the single node, say $u$, that has less than $p$ agents, actually no agents at this point of time, on it. Whenever the missing edge appears, $g_{i}$ divides into two groups, say $G_{1}$ and $G_{2}$, and starts executing DFS. As per Lemma 2, one of the groups $G_{1}$ or $G_{2}$ finds the node $u$ that has $<p$ agents and settles there. This continues until $u$ has at least $p$ settled agents. In the end, like our algorithm, when $G_{1}$ ($G_{2}$) is left with one agent, and it is sure that $G_{2}$ ($G_{1}$) has been dispersed, it settles at the current node, say $v$, as by this time the remaining $p$ agents of $g_{i}$ must have settled at $u$. Note that $v$ can be the node $u$ as well, but irrespective of what $v$ is, the dispersed configuration is achieved as exactly one node ($v$) contains $p+1$ agents.
  2. 2.
    Case $\bm{|g_{i}|=p}$ : In this case, $g_{j}$ with $p+1$ has been settled at some node $u_{1}$ in the earlier round. In observation 1, we have given the number of rounds of termination. In this case, agents in $g_{i}$ wait till that round. Agents in $g_{i}$ become group $G_{1}$, and the largest ID at $g_{j}$ becomes $G_{2}$. They start executing DFS as per our algorithm. As per Claim 1, if $G_{1}$ waits for consecutive $16n^{2}$ odd rounds, then $G_{2}$ visits each node at least once. In this case, $G_{2}$ and $G_{1}$ meet, and they become a group of $p+1$ agents, which is an earlier case. If $G_{1}$ does not wait for consecutive $16n^{2}$ odd rounds, then it is moving on the current DFS. And in at most $16n^{2}\times 4n^{2}$ odd rounds, it reaches node $u$ and settles. In at most $64n^{2}$ odd rounds, the agent in $G_{2}$ becomes settled at its current node. This ensures that the agents achieve a dispersed configuration because, after $64n^{2}$ odd rounds, if $G_{2}$ has not met $G_{1}$, it indicates that $G_{1}$ has completed its depth-first search (DFS).
Time Complexity of solving $(pn+1)$-Dispersion on TVG:
Now we are ready to state the final theorem.
Theorem 4.1
$(pn+1)$-Dispersion on TVG can be solved in $O(n^{4}\cdot max(\log n,\,\log p))$ rounds using $\Theta(\log n)$ memory per agent.
Proof
The correctness is justified in the above discussion. Since agents are using the algorithm mentioned in Section 3. Therefore, $\Theta(\log n)$ memory per agent is required. Since $n$ many $g_{i}$s are executing the algorithm, therefore, in $O(n^{4}\cdot\log n)$ rounds, one of the aforementioned cases is possible. In the first case, agents take $O(n^{4}\cdot\log p)$ extra rounds to achieve the dispersed configuration, and in the second case if both groups meet with each other within $64n^{2}$ odd rounds, then they convert into the first case, and take $O(n^{4}\cdot\log p)$ extra rounds to achieve the dispersed configuration. In the second case, if they do not meet with each other, then $g_{i}$ has reached node $u$ within $64n^{2}$, and the dispersion has been achieved. Therefore, the total time is $O(n^{4}\cdot\log n)+O(n^{4}\cdot\log p)=O(n^{4}\cdot max(\log n,\,\log p))$. This completes the proof. ∎
4.2 $k$-Dispersion on TVG
In Section 2, we have shown that to solve $(pn)$-Dispersion on TVG, agents need 1-hop visibility and global communication. In this section, we will extend the algorithm of [16] to solve $k$-Dispersion on TVG when agents are equipped with 1-hop visibility and global communication. Recall that their algorithm works in the weak dynamic graph model in comparison to TVG, i.e., the adversary can delete or add edges arbitrarily, maintaining the graph in each round connected. Let’s refer to the algorithm as $Algo\_weak\_Disp(k)$. First, we recall a few results from their work.
Theorem 4.2
[16] Given $k\leq n$ agents placed arbitrarily on the nodes of any $n$-node graph $\mathcal{G}_{r}$, where $\mathcal{G}_{r}$ is the graph at round $r\geq 0$. $Algo\_weak\_Disp(k)$ solves the dispersion in $\Theta(k)$ rounds with $\Theta(\log k)$ bits at each agent in the synchronous setting with global communication and 1-hop visibility. Also, all agents understand that dispersion has been achieved and terminated.
Based on $Algo\_weak\_Disp(k)$, the authors have the following lemma.
Lemma 5
Now, we are ready to provide how we can extend $Algo\_weak\_Disp(k)$ to solve $k$-Dispersion on TVG when $k$ agents in $G$ need 1-hop visibility and global communication. Note that this idea works even if agents are scattered, agents do not know $n$, and the adversary deletes more than 1 edge while maintaining $\mathcal{G}$ connected.
The agents run $Algo\_weak\_Disp(k)$. Using Lemma 5, we can observe that the number of nodes without any agents is reduced by at least 1 in each round. Therefore, after $n$ rounds, all nodes have at least 1 agent. At each round $r$, agents check whether their neighbour is occupied by some agent or not and share this information using global communication with all agents. At some round $r$, when agents do not get such information, agents can understand that each node has at least 1 agent. At every round $r^{\prime}>r$, at each node $v$, the largest ID agent becomes a leader, and it shares the information of the number of agents at node $v$ along with its ID. In this way, at every round $r^{\prime}$ with the help of global communication, agents can learn the value of $k$ using the number of agents at other nodes, and they can also learn the value of $n$ based on the number of distinct IDs they receive. Therefore, they do not have to remember $k$ and $n$, and in each round $r^{\prime}$, they can learn $k$ and $n$. This information allows agents to compute $\lceil k/n\rceil$. From round $r+2$, they start executing $Algo\_weak\_Disp(k)$. In each round, agents can learn $\lceil k/n\rceil$. This is important otherwise the memory requirement per agent will be $O(max(\log p,\log n))$. When they find at least one node has less than $\lceil k/n\rceil$ agents. They can follow $Algo\_weak\_Disp(k)$. Using Lemma 5, we can say that if there is a node $u$ with more than $\lceil k/n\rceil$ agents, then the extra agent from node $u$ moves to the other node $v$ which has less than $\lceil k/n\rceil$ agents. In this way, they can do this in each round. When agents find no node with more than $\lceil k/n\rceil$ agents, they can understand $k$-Dispersion on TVG has been achieved. Since $k$ agents are present, agents can achieve $k$-Dispersion on TVG in $O(k)$ rounds. In each round, agents need to remember their ID, and it does not have to remember $k$ or $n$. Therefore, agents need $O(\log n)$ memory as ID range is $[1,\,n^{c}]$, where $c$ is some constant. Using Lemma 1, we can say $\Theta(\log n)$ memory per agent is required. Based on this, we have the final theorem.
Theorem 4.3
[…]

### v3

@@ 7.2 Our algorithm with Correctness
      Else if $x\leq L$ and $y\geq l$, then it terminates.
    If $a_{i}$ is not the minimum ID agent at some node $u$, then it makes the following decision. Let $a_{l}$ be the minimum ID agent at node $u$. Since agent $a_{i}$ has the same information as agent $a_{l}$, it can calculate the decision of agent $a_{l}$ at round $r$. If it finds that agent $a_{l}$ decides to terminate in round $r$, then agent $a_{i}$ terminates. Otherwise, it does not do anything at round $r$.
Theorem 7.2.
Given $k\geq 1$ agents placed arbitrarily on the nodes of $G$. Our algorithm solves $k$-balanced dispersion in $O(k)$ rounds with $O(\log n)$ memory at each agent in the synchronous setting with global communication and 1-hop visibility. Also, all agents understand that dispersion has been achieved and terminated.
Proof.
If there is a hole and a multinode at round $r$, every agent receives this information. Due to Lemma 7.1, Algo_weak_disp(k)Algo\_weak\_disp(k) decreases the number of holes by at least one in every round. If $k\leq n$, then by Theorem 7.1, agents achieve dispersion within the first $n$ rounds and terminate. If $k>n$, initially there can be at most n1n-1 holes, so by Lemma 7.1, there will be no hole after the first $n$ rounds. After that, each agent $a_{i}$ sets $a_{i}.P=1$. Therefore, at any round $r$ where $a_{i}.P=1$, every node in $\mathcal{G}_{r}$ has at least one agent. Since each node is occupied, Algo_weak_disp(k)Algo\_weak\_disp(k) enables every agent to construct the map of $\mathcal{G}_{r}$. Therefore, in our algorithm at round $r$, graph $G^{\prime}$ is nothing but $\mathcal{G}_{r}$. Since each node has at least one agent, set $V_{1}$ is nothing but $V$, $k_{1}=k$ and $n_{1}=n$. Hence, agents correctly know that at most $L=\lceil k/n\rceil$ agents and at least $l=\lfloor k/n\rfloor$ agents can be present at a node. Once $a_{i}.P=1$, each agent $a_{i}$ can compute the same values of $x$, $y$, $U_{x}$, and $U_{y}$. Therefore, for every agent, the choice of the path $P$ between nodes $u_{1}$ and $w_{1}$ is unique in our algorithm. In this case, the size of $u_{1}$ decreases by 1 at the end of round $r$. Agents continue this process till they find that every node has at least $l$ agents and at most $L$ agents. Therefore, when agents terminate, agents are in a $k$-balanced dispersed configuration.
For $k\leq n$, dispersion is achieved in $O(k)$ rounds as per Theorem 7.1. For $k>n$, dispersion completes in $O(n+k)=O(k)$ rounds, where $O(n)$ to remove all holes, then $O(k)$ to maintain agents between $l$ and $L$ agents at every node, as only one node’s number of agents reduces per round. Agents only remember $a_{i}.P$ and $a_{i}.ID$, which requires $O(\log n)$ memory per agent. This completes the proof. ∎
Remark 7.1.
If $n+1\leq k\leq n^{c_{1}}$ for some constant $c_{1}$, then by Lemma 6.1, the memory required by the agents is $\Theta(\log n)$. Hence, in this range of $k$, the memory requirement by the agents is optimal.
8 Rooted $k$-balanced dispersion in 1-bounded 1-interval connected graphs
In Section 6, we have shown that solving $k$-balanced dispersion for $k=pn$, $p\in\mathbb{N}$, in 1-bounded 1-interval connected graphs requires both 1-hop visibility and 1-hop communication. In this section, we present two algorithms to solve $k$-balanced dispersion in 1-bounded 1-interval connected graphs when $k$ agents are co-located. The first algorithm is when $k=n+1$, and the second extends this to $k\in\{pn+1,pn+2\}$, where $p\in\mathbb{N}$. The question remains open whether 1-hop visibility and 1-hop communication are necessary to solve $k$-balanced dispersion for $k=pn+q$, where $q\in[3,n-1]$, in 1-bounded 1-interval connected graphs. We begin by describing the algorithm for $k=n+1$, and then generalise it to the case of $k=pn+1$ and $k=pn+2$. Both algorithms assume agents have face-to-face (f-2-f) communication, 0-hop visibility, and prior knowledge of $n$.
8.1 Dispersion with $n+1$ co-located agents
In the algorithm, we use the idea of a depth-first search (DFS) traversal by mobile agents from [2], which also solves the dispersion problem in static graphs $G$ for $k\leq n$ agents. Before describing the main idea of [2], we first introduce three technical terms used in their work, which we also use in our algorithm description.
  • 1.
    Settled agent: An agent $a_{i}$ is said to be settled at node $v$ if it occupies $v$ at round $r$ and never leaves $v$ in any round after $r$.
  • 2.
    Unsettled agent: An agent is unsettled if it is not settled.
  • 3.
    Root node of the DFS: The initial node from which an agent starts the DFS traversal is called the root node of the DFS.
The idea of [2] to solve dispersion for $k\leq n$ co-located agents is as follows. At each node $v$, a group of agents chooses an unexplored port $p$ and traverses through it: (i) if it reaches a node $u$ and finds that there is no agent settled at $u$, then one agent settles at $u$, and stores port $p^{\prime}$ (through which it entered node $u$) as its parent port. Now, the unsettled agents try to move through port $(p^{\prime}+1)$ mod $\delta_{u}$ in the explore state. (ii) Else, if the group of agents reach a node which already has a settled agent, the agent returns to node $v$ in the backtrack state. If the unsettled agents reach node $v$ in the backtrack state, then the agents compute $q=(p+1)$ mod $\delta_{v}$. If port $q$ is unexplored, it moves through port $q$ in the explore state. Else, if port $q$ equals the parent port of $v$, then the agents move through port $q$ in the backtrack state. The DFS algorithm takes $4m$ rounds to run on any graph, where $m$ is the number of edges in that graph. The information required for the group of agents at each node to run this algorithm can be stored by the settled agent that is already present at that node.
Recall that, according to our model, an agent does not know if any associated edge with respect to its current position is missing or not. This makes the adversary powerful as it can precompute the agents’ decision at a node and accordingly can delete an associated edge at the start of that round. If the group of unsettled agents decides to move through port $p$ according to our algorithm in round $r$, the adversary can remove the edge corresponding to port $p$ in round $r$. Thus, preventing the group of agents from moving in round $r$.
[…]